"Question2
A \, periodic \,function\, f(x) \,with \,period \,2 π \,is \,defined\, by :\\
 f(x)= [ (x+π)/2 , \, -π< x<0, \\(x-π)/2 , \, 0<x<π]\\

Determine \,the \,Fourier \,series \,for \,the \,periodic\, function."

1
Expert's answer
2021-08-05T11:04:14-0400

"a_0=\dfrac{2}{\pi}\displaystyle\int_{-\pi}^{\pi}f(x)dx"

"=\dfrac{2}{\pi}\displaystyle\int_{-\pi}^{0}\dfrac{x+\pi}{2}dx+\dfrac{2}{\pi}\displaystyle\int_{0}^{\pi}\dfrac{x-\pi}{2}dx"

"=\dfrac{2}{\pi}[\dfrac{x^2}{4}+\dfrac{\pi x}{2}]\begin{matrix}
 0 \\
 -\pi & d
\end{matrix}+\dfrac{2}{\pi}[\dfrac{x^2}{4}-\dfrac{\pi x}{2}]\begin{matrix}
 \pi \\
 0
\end{matrix}"

"=\dfrac{\pi}{2}-\dfrac{\pi}{2}=0"



"a_n=\dfrac{1}{\pi}\displaystyle\int_{-\pi}^{\pi}f(x)\cos{nx}dx"

"=\dfrac{1}{\pi}\displaystyle\int_{-\pi}^{0}\dfrac{x+\pi}{2}\cos{nx}dx+\dfrac{1}{\pi}\displaystyle\int_{0}^{\pi}\dfrac{x-\pi}{2}\cos{nx}dx"

"=-\dfrac{\cos{\pi n}-1}{2\pi n^2}+\dfrac{\cos{\pi n}-1}{2\pi n^2}=0"



"b_n=\dfrac{1}{\pi}\displaystyle\int_{-\pi}^{\pi}f(x)\sin{nx}dx"

"=\dfrac{1}{\pi}\displaystyle\int_{-\pi}^{0}\dfrac{x+\pi}{2}\sin{nx}dx+\dfrac{1}{\pi}\displaystyle\int_{0}^{\pi}\dfrac{x-\pi}{2}\sin{nx}dx"

"=-\dfrac{1}{2n}-\dfrac{1}{2n}=-\dfrac{1}{n}"




"f(x)=-\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{n}\sin{nx}"